# On Context Utilization in Summarization with Large Language Models

### v4

@@ 4.1 How is information in the middle treated?
Previous experiments show that LLMs place more emphasis on the beginning and the end of their context. We now narrow down on how LLMs treat the middle. To remove the effect of spread of salient information, we perform two controlled experiments in multi-document summarization. This setup enables us to shuffle the order of the input, which is not realistic for the single-document setup as it would break coherence. We only consider data points with the same number $k$ of documents: $k=7$ documents for Multi-XScience ($n=329$), and $k=5$ documents for Multi-News ($n=219$).
@@
Table 3: Performance in multi-document summarization on Multi-XScience (7 documents) and Multi-News (5 documents) when infilling the middle of the context window with random documents. R-2 is ROUGE-2, BS refers to BERTScore. % is the mean relative performance compared to the baseline with all documents.
@@ 4.3 Is scaling context length really useful?
Experiments from Section 4.1 confirm that LLMs struggle to summarize information contained in the middle of their context window. This poses issues for long-input summarization: after the initial part with (usually) high saliency, important information becomes sparser, and at the same time LLMs processing capability weakens. To investigate this issue, we infer long-document summarization with length truncated at $m*2k$ tokens, varying $m$ from 1 to 6.444Our hardware does not allow us to exceed 12k tokens. We use our longest context LLMs {Xgen-7B, Mistral-7B, GPT-3.5} ; as well as two open-source LLMs extending Llama-2-7B context window with position interpolation (Chen et al., 2023), a method gaining traction as an efficient way to scale LLMs’ context window. We use Vicuna-7B-1.5-16k [^5] In HuggingFace: lmsys/vicuna-7b-v1.5-16k, and Llama-2-7B-32k [^6] In HuggingFace: togethercomputer/LLaMA-2-7B-32K, with context of 16k tokens and 32k tokens, respectively.
@@ 6 Conclusion
Our work only considers open-source LLMs for summary generation and ignores closed-source LLMs such as the popular OpenAI’s GPT-3.5 and GPT-4, or Anthropic’s Claude. We made this decision to advocate for openness in LLM research ; yet we acknowledge that it would be interesting to also investigate properties of summaries generated by these paying API LLMs.
@@
Lastly, we can only evaluate a finite number of LLMs, and we settled for the evaluation of 5 recent and popular open-source LLMs. Findings may change as LLMs undergo changes and improvements in their training process.

### v6

@@ 4.1 How is information in the middle treated?
Previous experiments show that LLMs place more emphasis on the beginning and the end of their context. We now narrow down on how LLMs treat the middle. To remove the effect of spread of salient information, we perform two controlled experiments in multi-document summarization. This setup enables us to shuffle the order of the input, which is not realistic for the single-document setup as it would break coherence. We only consider data points with the same number $k$ of documents: $k=7$ documents for Multi-XScience ($n=329$), and $k=5$ documents for Multi-News ($n=219$) [^4] We don’t subsample from these subsets for GPT-3.5..
@@
Table 3: Performance in multi-document summarization on Multi-XScience (7 documents) and Multi-News (5 documents) when infilling the middle of the context window with random documents. R-2 is ROUGE-2, BS refers to BERTScore. % is the mean relative performance across all metrics compared to the baseline with all documents.
@@ 4.3 Is scaling context length really useful?
Experiments from Section 4.1 confirm that LLMs struggle to summarize information contained in the middle of their context window. This poses issues for long-input summarization: after the initial part with (usually) high saliency, important information becomes sparser, and at the same time LLMs processing capability weakens. To investigate this issue, we infer long-document summarization with length truncated at $m*2k$ tokens, varying $m$ from 1 to 6.555Our hardware does not allow us to exceed 12k tokens. We use our longest context LLMs {Xgen-7B, Mistral-7B, GPT-3.5} ; as well as two open-source LLMs extending Llama-2-7B context window with position interpolation (Chen et al., 2023), a method gaining traction as an efficient way to scale LLMs’ context window. We use Vicuna-7B-1.5-16k [^6] In HuggingFace: lmsys/vicuna-7b-v1.5-16k, and Llama-2-7B-32k [^7] In HuggingFace: togethercomputer/LLaMA-2-7B-32K, with context of 16k tokens and 32k tokens, respectively.
@@ 6 Conclusion
Our work mainly considers open-source LLMs for summary generation and ignores some popular closed-source LLMs such as Anthropic’s Claude. We made this decision to advocate for openness in LLM research ; yet we acknowledge that it would be interesting to also investigate properties of summaries generated by more paying API LLMs.
@@
Lastly, we can only evaluate a finite number of LLMs, and we settled for the evaluation of 6 recent and popular LLMs. Findings may change as LLMs undergo changes and improvements in their pre-training and instruction-tuning methodologies.